OpenStudy (anonymous):

Solve by Completing the square. b^(2)+2b+10=0

OpenStudy (anonymous):

b= -1 plus or minus 3i

OpenStudy (ash2326):

We know \[(a+b)^2=a^2+2ab+b^2\] here we have \[b^2+2b+10=0\] \[b^2+2b+1+9=0\] \[(b+1)^2+9=0\] or \[(b+1)^2=-9\] we get \[(b+1)=\pm 3i \] \[( i =\sqrt{-1} )\] we get \[b=-1\pm3i\]

OpenStudy (anonymous):

b^2+2b+10=0 b^2 + 1 = -10 +1 ( b + 1)^2 = -9 -> b+1 = +/- sqrt ( -9) = +/-3i => b = -1 +/-3i